A stone of mass 300g is thrown vertically upward at 16cm/s.find the maximum kinetic energy
1
Expert's answer
2020-09-21T06:28:07-0400

At the very beginning of the motion the potential enery was equal to 0 and, thus, the kinetic energy was maximum. During the motion the kinetic energy was decreasing up to 0 at the very top of the trajectory. After this it begun increasing up to the initial value (because of the energy conservation law) at the point of landing. Thus, a stone had maximum kinetic energy at the very beginning of the motion. As far as its velocity at this time was "v = 16cm\/s = 0.16m\/s", its kinetic energy was:


"K = \dfrac{mv^2}{2} = \dfrac{0.3kg\cdot (0.16m/s)^2}{2} = 0.00384J"

Answer. 0.00384 J.